OpenStudy (anonymous):

23,400 X 4 1/2% x 3 months =

OpenStudy (anonymous):

is this x or multiply sign

OpenStudy (anonymous):

multiple sign

OpenStudy (anonymous):

multiply sign'

OpenStudy (anonymous):

959400/(2/100 x3)= 959400/0.06=15990000

OpenStudy (anonymous):

i figured it out all ready. thanks anyway